Session: Feb 2002 (session 1)
Wavelength: 18. cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Sergio Mariotti at Tue Feb 19 10:58:26 MET 2002) Success
Observation is starded 10 minutes later

Torun (K.Borkowski at Tue Feb 19 11:05:38 MET 2002) Success
No known problems (remarks for this wavelength do not apply).

Effelsberg (A. Kraus at Tue Feb 19 13:33:17 MET 2002) Success

Westerbork (Tony Foley at Tue Feb 19 14:42:42 MET 2002) Minor failures
Poor phasing in 3 telescopes (RT5, RTC and RTD) gave SEFD increase of ~50% scans at 22:50 and 00:06 out of HA range.

Jodrell_Bank (Alastair Gunn at Fri Feb 22 10:08:12 MET 2002) Minor failures
Telescope parked due to high winds between 23:34 and 07:39 UT. No other known problems.

Cambridge (Alastair Gunn at Fri Feb 22 10:08:41 MET 2002) Failure
Cambridge did not observe due to serious fault with recorder.

Onsala (John Conway at Tue Feb 26 17:04:34 MET 2002) Failure
Did not observe due to failure of capstan motor. No tapes will be shipped